February: Baseball team collects season ticket revenue ($1200) for upcoming season. Season will run April-September. Monthly games are as follows: April 12 games, May 14, June 13, July 15,Aug 12,Sept 14.

How is the Initial cash collection ($1,200) recorded? How will the team recognize revenue? What do the revenue recognition say about this?
Answer #1

Total games to be held = 12 + 14 + 13 + 15 + 12 + 14 = 80 game

Income per game = $1,200 / 80 = $15

Initial cash recording

General Journal Debit Credit
Cash $1,200 -
Unearned revenues - $1,200

When even the game is held, income should be recognised. In April 12 games are held. Income per game is $15. Total income to be recognised in April is $15 X 12 games = $180. Out of $1,200 received, $180 should be recognised as income and remaining $1,020 (ie $1,200 - $180) should be treated as unrealised income.

Revenue recognitions says that when the games are conducted, then only income should be recognised in the books of accounts.